The phone company Ringular has a monthly plan where a customer pays a flat monthly fee, and then a set rate per minute. If a customer uses 110 minutes, the monthly cost will be $84. If the customer uses 960 minutes, the monthly cost will be $424.
Find the equation of the plan in the form y=mx+b
xis the number of monthly minutes used and
yis the total monthly cost of the Ringular plan.

Answer by Theo(13342)

at 110 minutes, the monthly cost is 84 dollars.
at 960 minutes the monthly cost is 424 dollars.

straight line equation is y = mx + b.
m is the slope.
b is the y-intercept which is the value of y when x = 0.

you have two points to work with.
in (x,y) format, they are (110,84) and (960,424).
x represents the number of minutes.
y represents the total cost for that number of minutes.

m is the slope which is equal to (y1-y1)/(x2-x1).
let (110,84) = (x1,y1) and let (960,424) = (x2,y2).
m = (424-84)/(960-110) = 340/850.
simplify to get 2/5.

y = mx + b becomes y = 2/5 * x + b
to find b, replace x and y with the values from one of the points.
any one of the points will work.
i chose (x2,y2) = (960,424)
y = 2/5 * x + b becomes 424 = 2/5 * 960 + b
solve for b to get:
b = 40

y = mx + b becomes y = 2/5 * x + 40

test with the 2 known points.
with x = 110, the equation becomes y = 2/5 * 110 + 40 = 84
with x = 960, the equation becomes y = 2/5 * 960 + 40 = 424

the equation you are looking for is confirmed to be y = 2/5 * x + 40
x is the number of minutes consumed per month.
y is the total cost per month.
there is a flat monthly fee of 40 dollars.
the cost per minute is 2/5 = .40 dollars = 40 cents.